# Arbitrary Precision Printed Ternary Neural Networks with Holistic Evolutionary Approximation

### v1

@@ V-C Comparison with the state of the art
Table II compares our approximate TNNs with the current state-of-the-art digital printed neural networks. We specifically evaluate against the approximate TNNs in [20], which use 11-bit inputs and require only ABCs, the approximate MLPs from [17], which employ a conservative approximate 88-bit multiplication scheme, and the approximate MLPs from [19, 18], which apply aggressive approximation using power-of-2 weights.  [17, 19, 18] also approximate the additions and use 44-bit inputs to enable high accuracy, despite the applied approximations. In Table II, we consider two accuracy loss thresholds relative to the baseline exact printed MLP [14]: $2\%$ for high accuracy and $5\%$ for high hardware gains. For improved readability, for each dataset and accuracy threshold, we report only the most area-efficient approximate printed MLP [17, 19, 18]. The table also includes the input precision used by each neural network, emphasizing the significance of our framework’s support for arbitrary input precision. As shown in Table II, different datasets and accuracy thresholds mandate varying precision for our TNNs. The reported area and power values include the analog-to-digital interface cost, which our approach minimizes by selecting the most area-efficient solution with respect to both digital classifier and interfacing costs. This often involves choosing the minimum necessary precision due to high ADC overheads. For consistency, the interfacing costs reported in Section III-A are used in all designs in Table II. Hereafter, references to ADC or interfacing costs refer to both ADC and/or ABC.
@@
The comparison with our prior work on approximate TNNs [20] further highlights the effectiveness of this solution. For 11-bit inputs (as in [20]) and identical accuracy, our TNNs achieve, on average, 1.81.8x and 2.92.9x lower area and power (including ADC costs), and 5.55.5x and 5.25.2x savings when considering only the TNN classifier. These gains (especially the latter) underline the superiority of our approximation strategy. [20] relies on approximate popcount units in both hidden and output layers, requiring multi-bit exact comparators in the hidden layer. In contrast, we fully approximate each hidden neuron as a single function, improving area-accuracy trade-offs for both the hidden neurons and the TNN overall. Moreover, [20] is restricted to 11-bit inputs due to its popcount-only approximation and uses only binary weights in the output layer for implementation simplicity. Therefore, [20] fails to meet the $2\%$ constraint for Cardio and WhiteWine, and does not even achieve a $5\%$ accuracy loss for Pendigits, Seeds, and Vertebral. Supporting arbitrary input precision and ternary weights in both layers, our framework delivers hardware-efficient solutions across all datasets and accuracy thresholds.
@@
Table II presents a comparative study against the printed neural networks state-of-the-art. For completeness, we also measure the accuracy that could be achieved by a silicon system, e.g., a deep MLP (more than 55 hidden layers and more than $50$ neurons per layer) that could run on a TinyML microcontroller. We observed that our TNNs remain in most cases within $5\%$ accuracy loss of the respective FP32 silicon-oriented MLP. The maximum accuracy drop is observed for the Arrhythmia dataset (i.e., 1616%), which, due to its increased number of input features, poses a priori a significant challenge for printed implementations [14]. From this analysis, we conclude that, given the immense computational limitations of printed technologies, our TNNs deliver reasonable accuracy compared to the silicon-oriented MLPs, while maintaining realistic area requirements and power consumption compatible with existing printed batteries—thus enabling feasible implementations. It is also noteworthy that these silicon-oriented MLPs require over 24,00024{,}000 MAC operations, which would translate to unrealistic area requirements in printed electronics.
Finally, we investigate the impact of our approximations on classification boundaries by analyzing the confidence margins of our approximate TNNs relative to the exact ones. By calculating the confidence margins for our approximate circuits of Table II, we observe that they are always maintained above 11 and often remain comparable to—or even better than—the exact TNN, demonstrating that classification boundaries are preserved under the applied approximations. Note, classification boundaries can be seamlessly incorporated into our optimization, by modifying (4) accordingly.
@@ V-E Execution Time Discussion
Training our arbitrary input precision TNNs took a maximum of $25$ minutes. After training, we identify the required LTG and popcount configurations; if any are missing from our library, we generate the corresponding approximate components. Fairly large LTGs are required with $n\cdot k$ reaching up to $89$. For each LTG approximation (i.e., threshold $\tau$), we perform three independent CGP runs, lasting between 3030min to 22h, based on the exact LTG’s complexity ($k\cdot n$). Similar timings are used to generate the approximate popcount units. However, all LTG and popcount approximations run concurrently, not impeding the scalability of our library generation. Finally, our multi-objective optimization, which approximates the TNN using our library of approximate components and NSGA-II, is very fast as it performs all evaluations in Python without requiring time-consuming hardware synthesis or simulations. In the worst case, it required only 66 minutes for Pendigits with 44-bit inputs. Experiments ran on a AMD Ryzen 5 3600 with 32GB RAM.

### v3

@@ V-C Comparison with the state of the art
Table II compares our approximate TNNs with the current state-of-the-art digital printed neural networks. We specifically evaluate against the approximate TNNs in [20], which use $1$-bit inputs and require only ABCs, the approximate MLPs from [17], which employ a conservative approximate $8$-bit multiplication scheme, and the approximate MLPs from [19, 18], which apply aggressive approximation using power-of-2 weights.  [17, 19, 18] also approximate the additions and use $4$-bit inputs to enable high accuracy, despite the applied approximations. In Table II, we consider two accuracy loss thresholds relative to the baseline exact printed MLP [14]: $2\%$ for high accuracy and $5\%$ for high hardware gains. For improved readability, for each dataset and accuracy threshold, we report only the most area-efficient approximate printed MLP [17, 19, 18]. The table also includes the input precision used by each neural network, emphasizing the significance of our framework’s support for arbitrary input precision. As shown in Table II, different datasets and accuracy thresholds mandate varying precision for our TNNs. The reported area and power values include the analog-to-digital interface cost, which our approach minimizes by selecting the most area-efficient solution with respect to both digital classifier and interfacing costs. This often involves choosing the minimum necessary precision due to high ADC overheads. For consistency, the interfacing costs reported in Section III-A are used in all designs in Table II. Hereafter, references to ADC or interfacing costs refer to both ADC and/or ABC.
@@
The comparison with our prior work on approximate TNNs [20] further highlights the effectiveness of this solution. For $1$-bit inputs (as in [20]) and identical accuracy, our TNNs achieve, on average, 1.81.8x and 2.92.9x lower area and power (including ADC costs), and 5.55.5x and 5.25.2x savings when considering only the TNN classifier. These gains (especially the latter) underline the superiority of our approximation strategy. [20] relies on approximate popcount units in both hidden and output layers, requiring multi-bit exact comparators in the hidden layer. In contrast, we fully approximate each hidden neuron as a single function, improving area-accuracy trade-offs for both the hidden neurons and the TNN overall. Moreover, [20] is restricted to $1$-bit inputs due to its popcount-only approximation and uses only binary weights in the output layer for implementation simplicity. Therefore, [20] fails to meet the $2\%$ constraint for Cardio and WhiteWine, and does not even achieve a $5\%$ accuracy loss for Pendigits, Seeds, and Vertebral. Supporting arbitrary input precision and ternary weights in both layers, our framework delivers hardware-efficient solutions across all datasets and accuracy thresholds.
@@
Table II presents a comparative study against the printed neural networks state-of-the-art. For completeness, we also measure the accuracy that could be achieved by a silicon system, e.g., a deep MLP (more than $5$ hidden layers and more than $50$ neurons per layer) that could run on a TinyML microcontroller. We observed that our TNNs remain in most cases within $5\%$ accuracy loss of the respective FP32 silicon-oriented MLP. The maximum accuracy drop is observed for the Arrhythmia dataset (i.e., 1616%), which, due to its increased number of input features, poses a priori a significant challenge for printed implementations [14]. From this analysis, we conclude that, given the immense computational limitations of printed technologies, our TNNs deliver reasonable accuracy compared to the silicon-oriented MLPs, while maintaining realistic area requirements and power consumption compatible with existing printed batteries—thus enabling feasible implementations. It is also noteworthy that these silicon-oriented MLPs require over 24,00024{,}000 MAC operations, which would translate to unrealistic area requirements in printed electronics.
Finally, we investigate the impact of our approximations on classification boundaries by analyzing the confidence margins of our approximate TNNs relative to the exact ones. By calculating the confidence margins for our approximate circuits of Table II, we observe that they are always maintained above $1$ and often remain comparable to—or even better than—the exact TNN, demonstrating that classification boundaries are preserved under the applied approximations. Note, classification boundaries can be seamlessly incorporated into our optimization, by modifying (4) accordingly.
@@ V-E Execution Time Discussion
Training our arbitrary input precision TNNs took a maximum of $25$ minutes. After training, we identify the required LTG and popcount configurations; if any are missing from our library, we generate the corresponding approximate components. Fairly large LTGs are required with $n\cdot k$ reaching up to $89$. For each LTG approximation (i.e., threshold $\tau$), we perform three independent CGP runs, lasting between 3030min to 22h, based on the exact LTG’s complexity ($k\cdot n$). Similar timings are used to generate the approximate popcount units. However, all LTG and popcount approximations run concurrently, not impeding the scalability of our library generation. Finally, our multi-objective optimization, which approximates the TNN using our library of approximate components and NSGA-II, is very fast as it performs all evaluations in Python without requiring time-consuming hardware synthesis or simulations. In the worst case, it required only $6$ minutes for Pendigits with $4$-bit inputs. Experiments ran on a AMD Ryzen 5 3600 with 32GB RAM.